# Data-driven controlled subgroup selection in clinical trials

### v1

@@ 3.1 Subgroup selection in the context of multivariate linear regression
One method due to wan2024confidence for conducting controlled subgroup selection in generalised linear models (GLMs) and mixed effects models is based on techniques for constructing simultaneous confidence bands in linear regression. Here, we briefly describe their procedure for the setting of GLMs.
@@
so that subgroup selection may be achieved by conducting inference on $x^{\top}\beta$ simultaneously over all values of $x\in\mathbb{R}^{d}$. For that purpose, let $\hat{\beta}$ be an estimate of $\beta$ based on independent copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, and assume that, asymptotically in $n$, $\sqrt{n}(\hat{\beta}-\beta)\sim\mathcal{N}_{d}(0,\Sigma)$. Further, let $\hat{\Sigma}$ be a consistent estimate of $\Sigma$ and let $\tilde{m}^{2}(x)$ be the estimated asymptotic variance of the linear part $\sqrt{n}\cdot x^{\top}\hat{\beta}$, that is, $\tilde{m}(x)=\sqrt{\widehat{\mathrm{Var}}\bigl(\sqrt{n}\cdot x^{\top}\hat{\beta}\bigr)}=\sqrt{x^{\top}\hat{\Sigma}x}$. For a given covariate region $K\subseteq\mathbb{R}^{d}$, wan2024confidence discuss the computation of critical values $\hat{c}_{1},\hat{c}_{2}\in[0,\infty)$ such that
@@
The amount by which the asymptotic coverage probability exceeds $1-\alpha$ depends on the data generating distribution, though the construction of $\hat{c}_{1}$ due to wan2024confidence ensures that the first two statements hold with equality for certain worst-case distributions.
In the special case of a normal linear model, i.e., where $g(t)=t$, so that $\eta(x)=x^{\top}\beta$, and $Y=X^{\top}\beta+\varepsilon$ for $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, independent of $X$, the stated Type I error control also holds for finite sample sizes when using ordinary least squares for estimating $\beta$ and $\sigma$. While the two-sided construction may be conservative in this setting, more recent work due to wan2025exact provides an alternative construction of $\hat{c}_{2}$, such that $1-\alpha$ is achieved exactly for certain distributions when $x$ consists of the intercept and one covariate.
@@ 3.2 Subgroup selection in the context of multivariate isotonic regression
mueller2024isotonic introduce an approach for subgroup selection in the context of multivariate isotonic regression based on martingale tests and multiple testing procedures for logically structured hypotheses. It is implemented in the R-package ISS (Mueller2023ISS_Code). This method has for instance been used to identify adults at risk of depression based on adverse childhood experiences (zhang2025acescorereplicablecombinations). Here we give a high-level summary of the method, beginning with the univariate case for simplicity of exposition.
@@
As before, let $(X,Y)$ denote a generic covariate-response-pair such that $\eta(x)=\mathbb{E}(Y\mid X=x)$ is increasing for $x\in\mathbb{R}$. If instead $\eta$ were decreasing, then one could consider the transformed covariate-response-pair (X,Y)(-X,Y) for which one would then have an increasing relationship. Hence, in everything that follows, one could replace the assumption of $\eta$ being increasing with the assumption that $\eta$ is monotone with known direction (either increasing or decreasing). Given access to independent copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, first fix $x_{0}$ to be the largest value among $X_{1},\ldots,X_{n}$. In order to construct a test of the null hypothesis***As shown by mueller2024isotonic, the described procedure will be valid for any realisation of $X_{1},\ldots,X_{n}$, so that we may consider them as fixed throughout. that $\eta(x_{0})<\tau$, one defines a test sequence, where each element $T_{k}$, $k=1,\ldots,\#\{i:X_{i}\leq x_{0}\}$, comprises the $k$ responses whose covariates are closest to $x_{0}$. The null hypothesis is rejected if for any $k$ the test statistic $T_{k}$ exceeds a critical threshold $c_{k}$ depending on $\alpha$. The threshold sequence $c_{k}$ can be chosen such that this procedure yields a provably valid test of the null hypothesis, as illustrated in Sections 3.2.1 and 3.2.2. The reason for choosing such a sequential procedure is that under the alternative hypothesis that $\eta(x_{0})\geq\tau$, the set of covariates $X_{1},\ldots,X_{n}$ having corresponding regression function values $\eta(X_{1}),\ldots,\eta(X_{n})$ at least at level $\tau$ is unknown. Supposing the hypothesis $\eta(x_{0})<\tau$ is rejected at level $\alpha$, the covariate-response pair with covariate value $x_{0}$ is removed from the data set, and the above testing procedure is repeated with $x_{0}$ set to the next-largest covariate value. If at any point the null hypothesis $\eta(x_{0})<\tau$ is not rejected at level $\alpha$ (or if all data points have been removed from the data set), then $x^{*}$ is taken to be the last value of $x_{0}$ for which the null was rejected and the procedure outputs $\hat{L}_{\alpha}:=[x^{*},\infty)$. If no null hypothesis is rejected, then $\hat{L}_{\alpha}$ will be the empty set. An upper confidence bound $\hat{U}_{\alpha}$ on $S_{\tau}$ can be obtained in a similar way, by calculating $\hat{L}_{\alpha}$ on transformed data $(-X_{1},-Y_{1}),\ldots,(-X_{n},-Y_{n})$ with a transformed threshold $\tau$, where the resulting set is denoted $\tilde{L}$, and then letting $\hat{U}_{\alpha}:=\mathbb{R}\setminus\{x\in\mathbb{R}:-x\in\tilde{L}\}$. Two-sided confidence sets can be obtained by letting $\hat{U}^{\mathrm{2S}}_{\alpha}:=\hat{U}_{\alpha/2}$ and $\hat{L}^{\mathrm{2S}}_{\alpha}:=\hat{L}_{\alpha/2}$.
@@
The procedure above consists of two nested loops: in the outer loop, the choice of $x_{0}$ is varied from the largest value in $X_{1},\ldots,X_{n}$ to the smallest, and the null hypothesis $\eta(x_{0})<\tau$ is formulated at each step, while in the inner loop, an associated test statistic is calculated by iterating through values of $k$. While the critical thresholds $c_{k}$ need to account for the multiplicity in the sequential structure of the test, no such multiplicity correction is present in the outer loop, where the test may be carried out at level $\alpha$ each time. The reason for this lies in the fact that the procedure terminates immediately if at any step the null hypothesis is not rejected, so that family-wise error rate control is guaranteed (hsu1999stepwise; westfall2001optimally).
@@
With additional work, a similar procedure can be developed for the multivariate setting. More specifically, since here $X_{1},\ldots,X_{n}$ are vectors in $\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, there is no natural total ordering: we may have two covariate vectors, neither of which is larger than the other in all components at once. Such vectors are said to be incomparable points or to be forming an antichain; this will play a role in Section 3.2.3, where categorical variables are discussed. For the construction of the test sequence, this issue may be remedied by considering only responses whose covariate vectors are smaller in every component than the vector $x_{0}\in\mathbb{R}^{d}$ and letting them enter the test sequence $T_{k}$ in the order of the covariate vectors’ supremum-norm distance from $x_{0}$. Designing an appropriate multiple testing procedure to constitute the outer loop described above is a little more involved, but mueller2024isotonic propose a method that builds on the idea of constructing a directed acyclic graph representing a logical structure between the null hypotheses to be tested. The $\alpha$-budget is then iteratively re-assigned to these hypotheses to again guarantee family-wise error rate control; see also work by bretz2009graphical, goeman2010sequential, and meijer2015multiple.
@@ 3.2.1 Binary responses
If the responses are binary, i.e. they take values 0 or $1$, then $Y_{i}$ given $X_{i}=x$ follows a Bernoulli distribution with the probability of observing a $1$ given by $\eta(x)$. In line with the general ideas outlined above, given a point $x_{0}=\bigl(x_{0}^{(1)},\ldots,x_{0}^{(d)}\bigr)\in\mathbb{R}^{d}$ at which we wish to test the null hypothesis $\eta(x_{0})<\tau$, we propose to order data points whose covariate value in each dimension is at most that of $x_{0}$ according to the supremum-norm distance from $x_{0}$. Denote these ordered covariate vectors by $X_{(1)},\ldots,X_{(n(x_{0}))}$, for $n(x_{0}):=\#\bigl\{i:X_{i}^{(j)}\leq x_{0}^{(j)}\text{ for all }j\in\{1,\ldots,d\}\bigr\}$, and let $Y_{(1)},\ldots,Y_{(n(x_{0}))}$ be the correspondingly ordered responses. Then the sequence $T_{k}:=\sum_{j=1}^{k}Y_{(j)}$ for $k\in\{1,\ldots,n(x_{0})\}$ tracks a cumulative count of the number of $1$s, starting with the points expected to be most indicative of the value of $\eta$ at $x_{0}$. Under the null hypothesis, every $Y_{(j)}$ takes the value $1$ with probability less than $\tau$, so $T_{k}$ should not grow to quickly (in particular, on average no faster than $k\tau$). This enables the construction of critical thresholds $c_{k}$ such that the null hypothesis should be rejected whenever $T_{k}>c_{k}$ for some $k$. mueller2024isotonic show that this idea may be formulated as a $p$-value, via the formula
@@ 3.2.2 Continuous responses
There are different ways in which continuous responses can be handled. For instance, mueller2024isotonic propose a way of constructing $p$-values targeting the specific case in which the responses may be assumed to follow a homoscedastic Gaussian distribution around their conditional expectation. An approach relying even less on distributional assumptions is based on the idea of considering quantiles of the response distribution. Here, the idea is to take the original responses $Y_{i}$ and dichotomize them into $Y^{*}_{i}:=\mathbbm{1}\{Y_{i}>\tau\}$. Thus, when applying the $p$-value for binary responses presented in Section 3.2.1 to the data $(X_{i},Y^{*}_{i}),i\in\{1,\ldots,n\}$, a point $x_{0}\in\mathbb{R}^{d}$, and a new threshold $\tau^{*}$, we yield a $p$-value for the null hypothesis that $\mathbb{E}(Y^{*}_{i}\mid X_{i}=x_{0})=\mathbb{P}(Y_{i}>\tau\mid X_{i}=x_{0})\leq\tau^{*}$. In particular, if we choose $\tau^{*}=1/2$, we are testing the null hypothesis that the conditional median of $Y_{i}$ (given $X_{i}=x_{0}$) is less than $\tau$. As a result, under the general assumption that the distribution of the responses is symmetric about their conditional expectation, this approach also leads to recovering $S_{\tau}$ with the same Type I error rate guarantees, in the setting of continuous responses.
As highlighted in mueller2024isotonic, further methods exist, e.g. in the setting of bounded or sub-Gaussian response distributions.
@@ 3.2.3 Categorical covariates
3.3 Subgroup selection for heterogeneous treatment effect
Meta-learners are advanced machine learning models that can be used to estimate the Conditional Average Treatment Effect (CATE), which measures the effect of a treatment on an outcome, conditional on covariates. The process involves an intermediate step where the meta-learner generates pseudo-outcomes. These pseudo-outcomes serve as surrogates or proxies for the individual treatment effects (ITE) similar to the “Virtual Twins” method due to foster2011subgroup, and thus provides a clearer picture of the treatment effect on an individual level. As a result, using the methods described above on top of the pseudo-outcomes can be a powerful approach for identifying subgroups where the treatment effects surpasses a predefined level.
We now introduce ITE and CATE more formally, before explaining how the discussed methods can be tailored to target these quantities. We will work in the setting of potential outcomes (hernan2025causal); that is, we observe a triple $(Y_{i},T_{i},X_{i})$ for each patient $i\in\{1,\ldots,n\}$, where $Y_{i}$ is a real-valued response, $T_{i}$ a binary treatment indicator taking values in $\{0,1\}$ and $X_{i}$ a $d$-dimensional, real covariate vector. Throughout, we assume these triples are independent and identically distributed between patients. In many settings, $T_{i}=0$ corresponds to patient $i$ having been assigned to a control arm, while $T_{i}=1$ signifies that they were assigned to a (novel) treatment. The potential outcomes framework then builds on the idea that we observe $Y_{i}=Y_{i}(1)$ when $T_{i}=1$ and $Y_{i}=Y_{i}(0)$ when $T_{i}=0$, but we never have access to the responses under treatment and control, $Y_{i}(1)$ and $Y_{i}(0)$ respectively, for the same patient at the same time. The individual treatment effect is the unobservable difference between these two quantities;

### v2

@@ 3.1 Subgroup selection in the context of multivariate linear regression
One method due to Wan et al. (2024) for conducting controlled subgroup selection in generalised linear models (GLMs) and mixed effects models is based on techniques for constructing simultaneous confidence bands in linear regression. Here, we briefly describe their procedure for the setting of GLMs.
@@
so that subgroup selection may be achieved by conducting inference on $x^{\top}\beta$ simultaneously over all values of $x\in\mathbb{R}^{d}$. For that purpose, let $\hat{\beta}$ be an estimate of $\beta$ based on independent copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, and assume that, asymptotically in $n$, $\sqrt{n}(\hat{\beta}-\beta)\sim\mathcal{N}_{d}(0,\Sigma)$. Further, let $\hat{\Sigma}$ be a consistent estimate of $\Sigma$ and let $\tilde{m}^{2}(x)$ be the estimated asymptotic variance of the linear part $\sqrt{n}\cdot x^{\top}\hat{\beta}$, that is, $\tilde{m}(x)=\sqrt{\widehat{\mathrm{Var}}\bigl(\sqrt{n}\cdot x^{\top}\hat{\beta}\bigr)}=\sqrt{x^{\top}\hat{\Sigma}x}$. For a given covariate region $K\subseteq\mathbb{R}^{d}$, Wan et al. (2024) discuss the computation of critical values $\hat{c}_{1},\hat{c}_{2}\in[0,\infty)$ such that
@@
The amount by which the asymptotic coverage probability exceeds $1-\alpha$ depends on the data generating distribution, though the construction of $\hat{c}_{1}$ due to Wan et al. (2024) ensures that the first two statements hold with equality for certain worst-case distributions.
In the special case of a normal linear model, i.e., where $g(t)=t$, so that $\eta(x)=x^{\top}\beta$, and $Y=X^{\top}\beta+\varepsilon$ for $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, independent of $X$, the stated Type I error control also holds for finite sample sizes when using ordinary least squares for estimating $\beta$ and $\sigma$. While the two-sided construction may be conservative in this setting, more recent work due to Wan et al. (2025) provides an alternative construction of $\hat{c}_{2}$, such that $1-\alpha$ is achieved exactly for certain distributions when $x$ consists of the intercept and one covariate.
@@ 3.2 Subgroup selection in the context of multivariate isotonic regression
Müller et al. (2025) introduce an approach for subgroup selection in the context of multivariate isotonic regression based on martingale tests and multiple testing procedures for logically structured hypotheses. It is implemented in the R-package ISS (Müller et al., 2023). This method has for instance been used to identify adults at risk of depression based on adverse childhood experiences (Zhang et al., 2025). Here we give a high-level summary of the method, beginning with the univariate case for simplicity of exposition.
@@
As before, let $(X,Y)$ denote a generic covariate-response-pair such that $\eta(x)=\mathbb{E}(Y\mid X=x)$ is increasing for $x\in\mathbb{R}$. If instead $\eta$ were decreasing, then one could consider the transformed covariate-response-pair (X,Y)(-X,Y) for which one would then have an increasing relationship. Hence, in everything that follows, one could replace the assumption of $\eta$ being increasing with the assumption that $\eta$ is monotone with known direction (either increasing or decreasing). Given access to independent copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, first fix $x_{0}$ to be the largest value among $X_{1},\ldots,X_{n}$. In order to construct a test of the null hypothesis***As shown by Müller et al. (2025), the described procedure will be valid for any realisation of $X_{1},\ldots,X_{n}$, so that we may consider them as fixed throughout. that $\eta(x_{0})<\tau$, one defines a test sequence, where each element $T_{k}$, $k=1,\ldots,\#\{i:X_{i}\leq x_{0}\}$, comprises the $k$ responses whose covariates are closest to $x_{0}$. The null hypothesis is rejected if for any $k$ the test statistic $T_{k}$ exceeds a critical threshold $c_{k}$ depending on $\alpha$. The threshold sequence $c_{k}$ can be chosen such that this procedure yields a provably valid test of the null hypothesis, as illustrated in Sections 3.2.1 and 3.2.2. The reason for choosing such a sequential procedure is that under the alternative hypothesis that $\eta(x_{0})\geq\tau$, the set of covariates $X_{1},\ldots,X_{n}$ having corresponding regression function values $\eta(X_{1}),\ldots,\eta(X_{n})$ at least at level $\tau$ is unknown. Supposing the hypothesis $\eta(x_{0})<\tau$ is rejected at level $\alpha$, the covariate-response pair with covariate value $x_{0}$ is removed from the data set, and the above testing procedure is repeated with $x_{0}$ set to the next-largest covariate value. If at any point the null hypothesis $\eta(x_{0})<\tau$ is not rejected at level $\alpha$ (or if all data points have been removed from the data set), then $x^{*}$ is taken to be the last value of $x_{0}$ for which the null was rejected and the procedure outputs $\hat{L}_{\alpha}:=[x^{*},\infty)$. If no null hypothesis is rejected, then $\hat{L}_{\alpha}$ will be the empty set. An upper confidence bound $\hat{U}_{\alpha}$ on $S_{\tau}$ can be obtained in a similar way, by calculating $\hat{L}_{\alpha}$ on transformed data $(-X_{1},-Y_{1}),\ldots,(-X_{n},-Y_{n})$ with a transformed threshold $\tau$, where the resulting set is denoted $\tilde{L}$, and then letting $\hat{U}_{\alpha}:=\mathbb{R}\setminus\{x\in\mathbb{R}:-x\in\tilde{L}\}$. Two-sided confidence sets can be obtained by letting $\hat{U}^{\mathrm{2S}}_{\alpha}:=\hat{U}_{\alpha/2}$ and $\hat{L}^{\mathrm{2S}}_{\alpha}:=\hat{L}_{\alpha/2}$.
@@
The procedure above consists of two nested loops: in the outer loop, the choice of $x_{0}$ is varied from the largest value in $X_{1},\ldots,X_{n}$ to the smallest, and the null hypothesis $\eta(x_{0})<\tau$ is formulated at each step, while in the inner loop, an associated test statistic is calculated by iterating through values of $k$. While the critical thresholds $c_{k}$ need to account for the multiplicity in the sequential structure of the test, no such multiplicity correction is present in the outer loop, where the test may be carried out at level $\alpha$ each time. The reason for this lies in the fact that the procedure terminates immediately if at any step the null hypothesis is not rejected, so that family-wise error rate control is guaranteed (Hsu and Berger, 1999; Westfall and Krishen, 2001).
@@
With additional work, a similar procedure can be developed for the multivariate setting. More specifically, since here $X_{1},\ldots,X_{n}$ are vectors in $\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, there is no natural total ordering: we may have two covariate vectors, neither of which is larger than the other in all components at once. Such vectors are said to be incomparable points or to be forming an antichain; this will play a role in Section 3.2.3, where categorical variables are discussed. For the construction of the test sequence, this issue may be remedied by considering only responses whose covariate vectors are smaller in every component than the vector $x_{0}\in\mathbb{R}^{d}$ and letting them enter the test sequence $T_{k}$ in the order of the covariate vectors’ supremum-norm distance from $x_{0}$. Designing an appropriate multiple testing procedure to constitute the outer loop described above is a little more involved, but Müller et al. (2025) propose a method that builds on the idea of constructing a directed acyclic graph representing a logical structure between the null hypotheses to be tested. The $\alpha$-budget is then iteratively re-assigned to these hypotheses to again guarantee family-wise error rate control; see also work by Bretz et al. (2009), Goeman and Solari (2010), and Meijer and Goeman (2015).
@@ 3.2.1 Binary responses
If the responses are binary, i.e. they take values 0 or $1$, then $Y_{i}$ given $X_{i}=x$ follows a Bernoulli distribution with the probability of observing a $1$ given by $\eta(x)$. In line with the general ideas outlined above, given a point $x_{0}=\bigl(x_{0}^{(1)},\ldots,x_{0}^{(d)}\bigr)\in\mathbb{R}^{d}$ at which we wish to test the null hypothesis $\eta(x_{0})<\tau$, we propose to order data points whose covariate value in each dimension is at most that of $x_{0}$ according to the supremum-norm distance from $x_{0}$. Denote these ordered covariate vectors by $X_{(1)},\ldots,X_{(n(x_{0}))}$, for $n(x_{0}):=\#\bigl\{i:X_{i}^{(j)}\leq x_{0}^{(j)}\text{ for all }j\in\{1,\ldots,d\}\bigr\}$, and let $Y_{(1)},\ldots,Y_{(n(x_{0}))}$ be the correspondingly ordered responses. Then the sequence $T_{k}:=\sum_{j=1}^{k}Y_{(j)}$ for $k\in\{1,\ldots,n(x_{0})\}$ tracks a cumulative count of the number of $1$s, starting with the points expected to be most indicative of the value of $\eta$ at $x_{0}$. Under the null hypothesis, every $Y_{(j)}$ takes the value $1$ with probability less than $\tau$, so $T_{k}$ should not grow to quickly (in particular, on average no faster than $k\tau$). This enables the construction of critical thresholds $c_{k}$ such that the null hypothesis should be rejected whenever $T_{k}>c_{k}$ for some $k$. Müller et al. (2025) show that this idea may be formulated as a $p$-value, via the formula
@@ 3.2.2 Continuous responses
There are different ways in which continuous responses can be handled. For instance, Müller et al. (2025) propose a way of constructing $p$-values targeting the specific case in which the responses may be assumed to follow a homoscedastic Gaussian distribution around their conditional expectation. An approach relying even less on distributional assumptions is based on the idea of considering quantiles of the response distribution. Here, the idea is to take the original responses $Y_{i}$ and dichotomize them into $Y^{*}_{i}:=\mathbbm{1}\{Y_{i}>\tau\}$. Thus, when applying the $p$-value for binary responses presented in Section 3.2.1 to the data $(X_{i},Y^{*}_{i}),i\in\{1,\ldots,n\}$, a point $x_{0}\in\mathbb{R}^{d}$, and a new threshold $\tau^{*}$, we yield a $p$-value for the null hypothesis that $\mathbb{E}(Y^{*}_{i}\mid X_{i}=x_{0})=\mathbb{P}(Y_{i}>\tau\mid X_{i}=x_{0})\leq\tau^{*}$. In particular, if we choose $\tau^{*}=1/2$, we are testing the null hypothesis that the conditional median of $Y_{i}$ (given $X_{i}=x_{0}$) is less than $\tau$. As a result, under the general assumption that the distribution of the responses is symmetric about their conditional expectation, this approach also leads to recovering $S_{\tau}$ with the same Type I error rate guarantees, in the setting of continuous responses.
As highlighted in Müller et al. (2025), further methods exist, e.g. in the setting of bounded or sub-Gaussian response distributions.
@@ 3.2.3 Categorical covariates
3.3 Subgroup selection for heterogeneous treatment effects
Meta-learners are advanced machine learning models that can be used to estimate the Conditional Average Treatment Effect (CATE), which measures the effect of a treatment on an outcome, conditional on covariates (Curth and van der Schaar, 2021; Kennedy, 2023; Lipkovich et al., 2024). In this section, we present one approach to using the meta-learner framework in the context of controlled subgroup selection.
First, we introduce individual treatment effects (ITE) and CATE formally. We will work in the setting of potential outcomes (Hernan and Robins, 2025); that is, we observe a triple $(Y_{i},T_{i},X_{i})$ for each patient $i\in\{1,\ldots,n\}$, where $Y_{i}$ is a real-valued response, $T_{i}$ a binary treatment indicator taking values in $\{0,1\}$ and $X_{i}$ a $d$-dimensional, real covariate vector. Throughout, we assume these triples are independent and identically distributed between patients. In many settings, $T_{i}=0$ corresponds to patient $i$ having been assigned to a control arm, while $T_{i}=1$ signifies that they were assigned to a (novel) treatment. The potential outcomes framework then builds on the idea that we observe $Y_{i}=Y_{i}(1)$ when $T_{i}=1$ and $Y_{i}=Y_{i}(0)$ when $T_{i}=0$, but we never have access to the responses under treatment and control, $Y_{i}(1)$ and $Y_{i}(0)$ respectively, for the same patient at the same time. The individual treatment effect is the unobservable difference between these two quantities;